OpenStudy (s3a):

Problem: A curve given parametrically by (x, y, z) = (2 + 3t, 2 – 2t^2, -3t – 2t^3). There is a unique point P on the curve with the property that the tangent line at P passes through the point (-10, -22, 76). Answer: P = (-4, -6, 22) I would appreciate it if someone could please tell me how to arrive at that answer!

OpenStudy (turingtest):

are you familiar with the formula for a parametric line\[\vec r(t)=\langle x_0,y_0,z_0\rangle+t\langle a,b,c\rangle\]?

OpenStudy (s3a):

I had forgotten about it, but you refreshed my mind, I suppose. :)

OpenStudy (turingtest):

\(\langle a,b,c\rangle\) is a vector along the line, and \(\langle x_0,y_0,z_0\rangle\) is a point on the line. hopefully you can see how this traces out a line in space parametrically as t increases

OpenStudy (turingtest):

for more detail: http://tutorial.math.lamar.edu/Classes/CalcII/EqnsOfLines.aspx

OpenStudy (turingtest):

in your problem, you have to find the initial point, given the line. from the equation of the line we can get our vector that points along the line tangentially by taking the derivative

OpenStudy (turingtest):

\[\vec r'(t)=\langle 3,-4t,-3-6t^2\rangle\]which in this case is your \[\langle a,b,c\rangle\]

OpenStudy (s3a):

(-10, -22, 76) = (x_0, y_0, z_0)?

OpenStudy (s3a):

Actually, wait.

OpenStudy (s3a):

(x_0, y_0, z_0) is the y-intercept of the regular line, and (-10, -22, 76) is the y-intercept of the tangent line, right?

OpenStudy (s3a):

Or, rather, the "anchor".

OpenStudy (s3a):

Not intercept, right?

OpenStudy (turingtest):

nor sure why you think this has to do with the y-intercept your \(\langle x_0,y_0,z_0\rangle\) is on your given parametric line...

OpenStudy (turingtest):

\(\langle x_0,y_0,z_0\rangle\) is actually what we're looking for since we know it's on the line we can write\[\langle x_0,y_0,z_0\rangle+t\langle a,b,c\rangle=\vec r(t)+t\vec r'(t)\]

OpenStudy (turingtest):

since we know that this line goes through point P for some t, we then can write\[\vec P=\vec r(t)+t\vec r'(t)\]

OpenStudy (s3a):

Is <x_o, y_0, z_0> a point that's on r(t) and r'(t) or just r'(t)?

OpenStudy (turingtest):

there is no such thing as a point "on" r'(t), r'(t) is a function that describes the way the tangent arrow on the curve r(t) faces at any given time. it cannot be said to "trace out a curve" like r(t) can. so the answer is that <x_0, y_0, z_0> is a point that's on r(t)

OpenStudy (s3a):

Ok, that makes sense, but let me think about the other stuff you said. 2 secs.

OpenStudy (turingtest):

sure thing :)

OpenStudy (turingtest):

take your time, I'm multitasking :P

OpenStudy (s3a):

Alright, lol. :)

OpenStudy (s3a):

So <a,b,c> = <c_1, c_2, c_2> where c_n is each component of dr/dt? Ok, so, I think I see that P = r + tr'. I'm going to do the work with my digital pen, and I'll get back to you. :)

OpenStudy (s3a):

Actually, wait.

OpenStudy (s3a):

Is r(t) = (x,y,z) or P + r(t) = (x,y,z)?

OpenStudy (s3a):

Wait, let me think more before I ask something, just in case it's a stupid question.

OpenStudy (s3a):

(I slept a little poorly today.)

OpenStudy (turingtest):

we didn't talk about x,y,z yet, but P is the point we are "pointing to", so your first statement was accurate

OpenStudy (s3a):

Is this "Is r(t) = (x,y,z)" what you're saying is accurate?

OpenStudy (turingtest):

such is college life :p <a,b,c> = <c_1, c_2, c_2> where c_n is each component of dr/dt - is correct

OpenStudy (s3a):

Oh, and, sadly, it is. :(

OpenStudy (turingtest):

r(t) is given, we want a specific point *on* r(t) that we called <x_0,y_0,z_0>

OpenStudy (s3a):

Okay, so, r(t) IS equal to (x,y,z)?

OpenStudy (s3a):

(Not saying (x_0,y_0,z_0).)

OpenStudy (turingtest):

I don't know what you really mean, but I'm going to say "no" let's say <x,y,z> is the line we are trying to find to be clear, call it something else, like v(t) r(t) is the curve we are given <x_0, y_0, z_0> is a point on that curve

OpenStudy (turingtest):

if it helps you to think of it this way, the equation of the line we want is\[\langle x_0,y_0,z_0\rangle+t\langle a,b,c\rangle=\vec r(t)+t\vec r'(t)=\langle x,y,z\rangle=\vec v(t)\]where v(t) is the line we need to find the equation of

OpenStudy (turingtest):

on the line v(t) is the point P somewhere, so we can say that there exist some x,y,z such that for some value of t, we have\[\vec v(t)=\vec P=\langle x,y,z\rangle=\langle-10,-22,76\rangle=\vec r(t)+t\vec r'(t)\]

OpenStudy (s3a):

Actually, you lost me with that. I can't relate that to r = P + tr'. I'm really sorry. :(

OpenStudy (turingtest):

solve for t

OpenStudy (s3a):

Wait, you seem to contradict yourself.

OpenStudy (s3a):

P = <x,y,z>? I thought P was <a,b,c>. <a,b,c> = <x,y,z>?

OpenStudy (s3a):

If P = <x,y,z> then r = P + tr' doesn't hold.

OpenStudy (turingtest):

no, <a,b,c> is r'(t) for some t

OpenStudy (turingtest):

it is the tangent vector we need to get our line

OpenStudy (s3a):

I meant P = <x_0, y_0, z_0>, I think.

OpenStudy (s3a):

Wait, I'm really sorry, I'm kind of a mess, and I'm trying really hard to focus! :@

OpenStudy (s3a):

Oh, wait!

OpenStudy (turingtest):

that's not true either, and no worries about it, my explanation may not be the best.. in general, the equation of a line parametrically is\[\vec v(t)=\langle x_0,y_0,z_0\rangle+t\langle a,b,c\rangle\]right?

OpenStudy (turingtest):

sorry my internet went out for a sec

OpenStudy (turingtest):

had a revelation?

OpenStudy (s3a):

I did, but I lost it kind of. Basically, what do we do after we have t?

OpenStudy (turingtest):

the question is asking for the point on the curve that has the tangent line we found, so we want to plug in the value of t into r(t)

OpenStudy (s3a):

Okay, so, I'm going to rewrite this all concisely, because of my sleep-deprived state when I scroll up and down too much, I forget what I was thinking.

OpenStudy (turingtest):

haha okay, let me try to write out a concise explanation in the meantime...

OpenStudy (s3a):

:$

OpenStudy (s3a):

Is r'(t) technically the derivative of r(t) - (x_0, y_0, z_0) which just happens to be the derivative of r(t)?

OpenStudy (s3a):

Wait, I take that back for now.

OpenStudy (turingtest):

ok lol

OpenStudy (s3a):

:D

OpenStudy (turingtest):

The equation of a line parametrically is\[\vec v(t)=\langle x,y,z\rangle=\langle x_0,y_0,z_0\rangle+t\langle a,b,c\rangle\]where \(\langle\ x_0,y_0,z_0\rangle\) is a point on the line, and \(\langle\ a,b,c\rangle\) is a vector in the same direction as the line \(\vec v(t)\). Note: do NOT confuse this with the equation of the given curve \(\vec r(t)\). I used the same variables at the beginning of my explanation and that probably did not help... In this case, \(\langle\ x_0,y_0,z_0\rangle\) is on the curve given, \(\vec r(t)\), and the line must be tangent to the curve, and so must point in the direction of \(\langle a,b,c\rangle=\vec r'(t)\) Putting all this together, this means we can write the equation of the tangent line to the curve we are looking for as\[\vec v(t)=\vec r(t)+t\vec r'(t)\] Now all we need to do is note that \(P\) is on \(\vec v(t)\) and we can write\[\vec P=\vec r(t)+t\vec r'(t)\]which can be solved for \(t\) Once you have that, you can get the point on the curve we started from, \(\langle x_0,y_0,z_0\rangle\), but plugging that value of \(t\) back into the equation of the curve \(\vec r(t)\), which gives you your answer.

OpenStudy (s3a):

Ok, so, (2 + 3t, 2 – 2t^2, -3t – 2t^3) = (-10, -22, 76) yields different values of t per component.

OpenStudy (turingtest):

you can't set the equation for the curve equal to the point; the point is not on the curve read what I wrote above and let me know where I lose you if you could :P

OpenStudy (s3a):

But, you said ⟨x,y,z⟩=⟨x0,y0,z0⟩+t⟨a,b,c⟩ = P I took P = (-10, -22, 76) and equated it to P = (x,y,z) = v(t).

OpenStudy (turingtest):

yeah that's right

OpenStudy (s3a):

(x,y,z) = (2 + 3t, 2 – 2t^2, -3t – 2t^3) so, (-10, -22, 76) = (2 + 3t, 2 – 2t^2, -3t – 2t^3), but that gives me different values of t per component.

OpenStudy (turingtest):

no, you are again confusing the equation of the line and that of the curve r(t)

OpenStudy (turingtest):

<x,y,z>=v(t) NOT r(t)

OpenStudy (s3a):

r(t) = (2 + 3t, 2 – 2t^2, -3t – 2t^3) ? if so, why does the problem say (x, y, z) = (2 + 3t, 2 – 2t^2, -3t – 2t^3)?

OpenStudy (turingtest):

<x,y,z>=r(t_0) + r'(t_0) for some t = t_0 as per my argument above

OpenStudy (s3a):

Okay, I think I should just do re-read this tomorrow. :P

OpenStudy (turingtest):

okay, terminology issue is all that is fine let's call the function r(t)=<x,y,z> like they do (which I didn't realize, sorry) then we need another name for the components in the line we are trying to find, so then we can call those v(t)=<p,q,r> the semantics are what's killing us here :P

OpenStudy (turingtest):

so let's say then that the line can be written v(t)=<p,q,r>, then we'll say v(t)=r(t)+t*r'(t) <p,q,r>=<x,y,z>+t<a,b,c> and for some particular value t=t_0, we can call the point <p_0,q_0,r_0>=P=<x_0,y_0,z_0>+t<a,b,c>

OpenStudy (turingtest):

change the line in my above explanation and say the general equation of a line is\[\vec v(t)=\langle p,q,r\rangle=\langle x_0,y_0,z_0\rangle+t\langle a,b,c\rangle\]if that helps...

OpenStudy (s3a):

Actually, I, honestly, think I'm too mentally drained to complete this problem successfully today, but I have a feeling that the last few posts will have cleared things up. I'll likely check it tomorrow, and, hopefully, report that I get it. :)

OpenStudy (s3a):

Thanks a lot for your explanations.

OpenStudy (turingtest):

sure thing, that often helps good luck, get some shut-eye!

OpenStudy (s3a):

I intend to. :D See you some other time. Bye for now!

OpenStudy (s3a):

@TuringTest: I changed the variables to be better compatible with my brain. :) I get the right answer, but did I get it correctly, or is it a fluke?: (x,y,z) = (2 + 3t, 2 - 2t^2, -3t - 2t^3) d/dt (x,y,z) = (3, -4t, -3, -6t^2) Let's call the point (-10,-22, 76) point Q.: Q = (-10,-22,76) Since we know that both the curve and one of its tangent lines pass through a point P and that that same tangent line also passes through the point we denoted by Q, the following must hold.: Q = (x,y,z) + t * d/dt (x,y,z) (-10,-22,76) = (2 + 3t, 2 - 2t^2, -3t - 2t^3) + t * (3, -4t, -3 - 6t^2) (-10,-22,76) = (2 + 3t + 3t, 2 - 2t^2 - 4t^2, -3t - 3t - 2t^3 - 6t^3) (-10,-22,76) = (2 + 6t, 2 - 6t^2, - 6t - 8t^3) From this we find the following three equations.: -10 = 2 + 6t ==> t = -12/6 = -2 -22 = 2 - 6t^2 ==> t = -2 holds for this equation too. 76 = -6t - 8t^3 ==> t = -2 holds for this equation too. Now that we know that t = -2, we can plug it back into (x,y,z) = (2 + 3t, 2 - 2t^2, -3t - 2t^3) to get (2 + 3*(-2), 2 - 2*(-2)^2, -3*(-2) - 2*(-2)^3) = (-4,-6,22).

OpenStudy (s3a):

I should note that when plugging back in -2 to (x,y,z), I'm plugging in back t_P = -2. (Emphasis on the subscript P of t_P.)

OpenStudy (turingtest):

yep, you got it :) just needed a little rest I guess

OpenStudy (s3a):

Thanks, again. :)